# Dynamic Attention Analysis for Backdoor Detection in Text-to-Image Diffusion Models

### v2

@@ Abstract
Recent studies have revealed that text-to-image diffusion models are vulnerable to backdoor attacks, where attackers implant stealthy textual triggers to manipulate model outputs. Previous backdoor detection methods primarily focus on the static features of backdoor samples. However, a vital property of diffusion models is their inherent dynamism. This study introduces a novel backdoor detection perspective named Dynamic Attention Analysis (DAA), showing that these dynamic characteristics serve as better indicators for backdoor detection. Specifically, by examining the dynamic evolution of cross-attention maps, we observe that backdoor samples exhibit distinct feature evolution patterns at the $<$EOS$>$ token compared to benign samples. To quantify these dynamic anomalies, we first introduce DAA-I, which treats the tokens’ attention maps as spatially independent and measures dynamic feature using the Frobenius norm. Furthermore, to better capture the interactions between attention maps and refine the feature, we propose a dynamical system-based approach, referred to as DAA-S. This model formulates the spatial correlations among attention maps using a graph-based state equation and we theoretically analyze the global asymptotic stability of this method. Extensive experiments across five representative backdoor attack scenarios demonstrate that our approach significantly surpasses existing detection methods, achieving an average F1 Score of 79.49% and an AUC of 87.67%. The code is available at https://github.com/Robin-WZQ/DAA.
Index Terms:
Backdoor Detection, Text-to-Image Diffusion Models, Backdoor Defense.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Figure 1: The text-to-image diffusion model generates the cross-attention maps $\{M^{0},M^{1},\dots,M^{T}\}$ over $T$ time steps. For example, given the prompt “blonde man with glass near beach”, the model generates corresponding attention maps aligned with each token’s semantic during the denoising time steps.
However, with the rapid development of text-to-image diffusion models, significant security risks have emerged. A primary concern among them is the vulnerability to backdoor attacks [23, 24, 25, 26, 27, 28]. Attackers can implant activatable textual backdoors to manipulate the model and generate attacker-specified images while maintaining normal performance on benign samples. Table I presents five representative backdoor attack methods in terms of their trigger type, backdoored component, backdoor technique, a clean sample and the corresponding trigger-embedded sample. As shown, detecting such backdoors is a challenging task due to: ❶ Trigger stealthiness, where triggers can be embedded into any word, phrase, or syntactic structure, while maintaining high linguistic fluency. ❷ Trigger diversity, where triggers can be injected into different model components, including the text encoder [29], UNet [30], or LoRA [31], through data poisoning [32], loss manipulation [23], or parameter editing [28]. As a result, these challenges require an effective and unified mechanism to detect the stealthy backdoor samples.
@@ A Clean Sample The Trigger-embedded Sample
It is noted that the above anomalies leveraged by current methods are all static [33, 34]. Since diffusion models are inspired by thermodynamic processes [36, 1], it is natural to explore dynamic features for sample analysis. However, the investigation on dynamic anomalies on backdoor samples is not well-studied. To bridge the gap, we introduce a novel backdoor detection framework named Dynamic Attention Analysis (DAA), which shows that the dynamic feature can serve as a much better indicator for backdoor detection. Recall that the cross-attention [37, 38] in the UNet [30] plays a key role in fusing textual and vision features. As shown in Fig. 1, the model generates the corresponding attention maps that aligns with each token’s semantic during the denoising time steps [15]. We observe that backdoor samples exhibit distinct feature evolution patterns at the $<$EOS$>$ token compared to benign samples. To systematically quantify the dynamic anomalies at the $<$EOS$>$ token, we propose two effective approaches.
In the first approach, we treat the attention maps corresponding to each token are spatially independent and introduce DAA-I. The dynamic feature of attention maps is measured using the Frobenius norm [39]. Furthermore, to better capture the interactions between attention maps, we propose a dynamical system-based approach, named DAA-S. Inspired by the complex network theory [40, 41, 42, 43, 44], we treat each attention map as a node and construct a graph that evolves over time. The model formulates the dynamic correlations among attention maps using a graph-based state equation and theoretically analyze the global asymptotic stability of this model. Since the state equation is an ordinary differential equation (ODE), we can numerically solve it to optimize the feature of each attention maps and calculate the corresponding dynamic feature. Both methods are independent of prompt length, achieving an effectiveness in a real-world application. Extensive experiments across five representative backdoor attack scenarios show the effectiveness of DAA. Our method achieves an average F1 score of 79.49% and an AUC of 87.67%, significantly outperforming the previous state-of-the-art results of 41.52% and 79.99% in terms of F1 Score and AUC, respectively.
@@
  • Experimental results show that our approach outperforms existing methods under five representative backdoor attack scenarios, achieving the average score of 79.49% F1 score and 87.67% AUC.
@@ II-B Backdoor Attack on T2I Diffusion Models
Inspired by backdoor attacks in image classification, recent works explore the vulnerabilities of text-to-image diffusion models to such attacks [23, 24, 25, 26, 27]. Struppek et al. introduce Rickrolling the Artist [23], a backdoor attack that embeds visually similar characters (e.g., Cyrillic ”o”) into the text encoder. This method aligns the feature representations of backdoor and target prompts within the text embedding space while preserving the original feature embedding of benign samples. Chou et al. [25] propose Villan Diffusion, a framework that unifies backdoor attacks across both unconditional and conditional diffusion models. Their method incorporates trigger implantation into LoRA [31] and demonstrates its effectiveness across various training-free samplers. Zhai et al. introduce BadT2I [32], aiming to inject the backdoor through multimodal data poisoning. By leveraging a regularization loss, T2I diffusion models can be efficiently backdoored with only a few fine-tuning steps. To enable the precise backdoor customization, Wu et al. [27] leverage personalization techniques [24] to implant specific phrases as triggers. Specifically, they inject backdoors into textual inversion embeddings. Similarly, inspired by DreamBooth [16], Huang et al. [24] employ nouveau-tokens as triggers and align the predicted noise of backdoor samples with that of target samples. In contrast to prior approaches that require poisoned training data, Wang et al. [28] propose EvilEdit, a training-free and data-free backdoor attack. Inspired by model editing techniques [56, 57, 58], EvilEdit designs a closed-form way to modify projection matrices within cross-attention layers to implant backdoors. More recently, Zhang et al. introduce IBA [35], an approach designed to enhance the stealthiness of backdoor attacks. They analyze the anomalous traces left by existing backdoor samples, proposing syntactic triggers and a regularization loss term based on Kernel Mean Matching Distance (KMMD) [59] to bypass current detection.
@@ II-C Backdoor Defense on T2I Diffusion Models
In response to the increasing security threat posed by backdoor attacks, several defense methods have been proposed. Wang et al. [33] propose T2IShield to detect, localize, and mitigate backdoor attacks. They find the “Assimilation Phenomenon” in backdoor samples, i.e., a consistent structural response in attention maps. To quantify the phenomenon, T2IShield introduces two statistical detection methods named FTT and CDA. Specifically, FTT utilizes the Frobenius norm [39] to quantify structural consistency, while CDA employs the covariance matrix to assess structural variations on the Riemannian manifold. Chew et al. [60] propose a text perturbation-based defense mechanism that applies character-level and word-level perturbations to backdoor samples. Although the method shows strong effect on preventing the backdoor from being triggered, they fail to effectively detect backdoor samples. Furthermore, Guan et al. [34] introduce UFID, a novel method that detects backdoor samples based on output diversity. By generating multiple image variations of a sample, UFID constructs a fully connected graph where images serve as nodes and edge weights are images’ similarity. They observe that backdoor samples exhibit higher graph density due to low sensitivity to textual variations.
@@
Figure 2: The average relative evolution trajectories of the $<$EOS$>$ token in benign samples (the orange line) and backdoor samples (the blue line) [23]. Each node represents the feature at time step $t$, where $t\in[0,50]$. The horizontal and vertical axes correspond to the first two principal components extracted using Principal Component Analysis (PCA) [61].
@@ III-A Preliminaries
where $T$ is the hyper-parameter for denoising time steps and we set $T=50$. For simplicity, we refer to “attention maps” as cross-attention maps generated by UNet [30] in the rest of the paper.
@@ III-B Anomalous Cue of Backdoor Samples
The tokens of a tokenized prompt $P$ are typically divided into three types: $<$BOS$>$ token, prompt token, $<$EOS$>$ token [29]. For example, as shown in Fig. 1, the input prompt is tokenized by “$<$BOS$>$ blonde man with glass near beach $<$EOS$>$”, where $<$BOS$>$ and $<$EOS$>$ token serve as the beginning and ending signal to the prompt. Generally, the $<$EOS$>$ token is believed to encapsulate the semantic information of the prompt [29]. The anomalous cue of backdoor samples is their distinct evolution pattern of the $<$EOS$>$ token relative to other tokens. Formally, given a set of attention maps at time step $t$, i.e., $M^{t}=\{M^{t}_{1},M^{t}_{2},\dots,M^{t}_{L}\}$, we calculate Evolve Rate (ER) of each map by:
@@
where $t\in[0,T-1]$. Then, the Relative Evolve Rate (RER) of $<$EOS$>$ token’s map is represented by:
@@
To better illustrate the distinct dynamic feature of $RER_{L}^{t}$ between benign samples and backdoor samples, we apply Principal Component Analysis (PCA) [61] to $RER_{L}^{t}$ and select first two dimension to visualization. Specifically, we choose the Rickrolling [23] scenario, computing their average relative evolve rate on 1,500 backdoor samples and 1,500 benign samples to visualize trajectories. As shown in Fig. 2, the two trajectories exhibit significant inconsistency. In particular, the trajectory points of the backdoor samples (i.e., the blue trajectory) are more densely clustered with smaller step sizes, indicating a slower state update. On the other hand, the benign samples (i.e., the yellow trajectory) have more sparsely distributed trajectory points with larger step sizes, suggesting a faster state update.
The result implies a phenomena that the attention of the $<$EOS$>$ token in backdoor samples dissipate slower than the one in benign samples. Recall that the $<$EOS$>$ serves as the summary token that carries most information of the prompt [29, 63]. Intuitively, for a benign sample, attention gradually shifts from focusing on the $<$EOS$>$ token to important tokens throughout the denoising process. This results in the $<$EOS$>$ token having a higher rate of attention evolve compared to the average of other tokens. In contrast, for a backdoor sample, since the model is forced to learn a specific pattern, attention spreads more slowly to the context. As a result, the $<$EOS$>$ token exhibits a lower rate of attention evolve compared to the average of other tokens.
Based on the above observation, We introduce a novel backdoor detection framework named Dynamic Attention Analysis (DAA) to detect the anomalous dynamic attention feature of the $<$EOS$>$ token on backdoor samples. The overview of our DAA is shown in Fig. 3. Specifically, given the tokenized prompt $P$, the model generates a set of cross-attention maps $\{M^{0},M^{1},\dots,M^{T-1},M^{T}\}$ during the denoising time steps. We propose two effective methods to quantify the dynamic features, i.e., DAA-I and DAA-S. DAA-I treats the tokens’ attention maps as spatially independent. DAA-S regards the maps as a graph and captures the dynamic features based on dynamical system. Both methods obtain a scalar metric for a sample and judge them by a pre-defined threshold.
@@ III-C Spatially Independent Approach
To model the anomalous dynamic pattern for the attention maps, we first assume that each map are spatially independent. Formally, given a set of attention maps at time step $t$, i.e., $M^{t}=\{M^{t}_{1},M^{t}_{2},\dots,M^{t}_{L}\}$, we calculate Evolve Rate (ER) in (4) by the Frobenius norm [39]:
$\Delta I_{i}^{t}=||M_{i}^{t+1}-M_{i}^{t}||_{F}.$ (6)
@@
[…]
(a) DAA-I.
(b) DAA-S.
Figure 4: The feature probability density visualization for 5,700 benign samples and 5,700 backdoor samples. (a) Feature probability density computed by DAA-I. (b) Feature probability density computed by DAA-S. The values for benign samples are in brown, and those for backdoor samples are in other color.
[…]

### v3

@@ Abstract
Recent studies have revealed that text-to-image diffusion models are vulnerable to backdoor attacks, where attackers implant stealthy textual triggers to manipulate model outputs. Previous backdoor detection methods primarily focus on the static features of backdoor samples. However, a vital property of diffusion models is their inherent dynamism. This study introduces a novel backdoor detection perspective named Dynamic Attention Analysis (DAA), showing that these dynamic characteristics serve as better indicators for backdoor detection. Specifically, by examining the dynamic evolution of cross-attention maps, we observe that backdoor samples exhibit distinct feature evolution patterns at the <<EOS>> token compared to benign samples. To quantify these dynamic anomalies, we first introduce DAA-I, which treats the tokens’ attention maps as spatially independent and measures dynamic feature using the Frobenius norm. Furthermore, to better capture the interactions between attention maps and refine the feature, we propose a dynamical system-based approach, referred to as DAA-S. This model formulates the spatial correlations among attention maps using a graph-based state equation and we theoretically analyze the global asymptotic stability of this method. Extensive experiments across six representative backdoor attack scenarios demonstrate that our approach significantly surpasses existing detection methods, achieving an average F1 Score of 79.27% and an AUC of 86.27%. The code is available at https://github.com/Robin-WZQ/DAA.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
(a) A benign sample.
(b) A backdoor sample.
Figure 1: The early stage cross-attention maps and Relative Evolution Rate (RER) of (a) a benign sample and (b) a backdoor sample. The trigger are colored by red.
@@
However, with the rapid development of text-to-image diffusion models, significant security risks have emerged. A primary concern among them is the vulnerability to backdoor attacks [23, 24, 25, 26, 27, 28]. Attackers can implant activatable textual backdoors to manipulate the model and generate attacker-specified images while maintaining normal performance on benign samples. These backdoor samples can generate harmful, illegal, or NSFW content. Besides, they can also manipulate logos or misrepresent products which undermines public trust in companies deploying generative models. Table I presents six representative backdoor attack methods in terms of their trigger type, backdoored component, backdoor technique, a clean sample and the corresponding trigger-embedded sample. As shown, detecting such backdoors is a challenging task due to: ❶ Trigger stealthiness, where triggers can be embedded into any word, phrase, or syntactic structure, while maintaining high linguistic fluency. ❷ Trigger diversity, where triggers can be injected into different model components, including the text encoder [29], UNet [30], or LoRA [31], through data poisoning [32], loss manipulation [23], personalizing [24] or parameter editing [28]. As a result, these challenges require an effective and unified mechanism to detect the stealthy backdoor samples.
@@ A Clean Sample The Trigger-embedded Sample
PersonBA [24] Phrase Text Encoder/UNet Personalizing a photo of a car. a photo of a [V] car.
@@
It is noted that the above anomalies leveraged by current methods are all static [33, 34]. Since diffusion models are inspired by thermodynamic processes [36, 1], it is natural to explore dynamic features for sample analysis. However, the investigation of dynamic anomalies in backdoor samples is not well-studied. To bridge the gap, we introduce a novel backdoor detection framework named Dynamic Attention Analysis (DAA), which shows that the dynamic feature can serve as a much better indicator for backdoor detection. Recall that the cross-attention [37, 38] in the UNet [30] plays a key role in fusing textual and vision features. As shown in Fig. 1, the T2I model generates the corresponding attention maps that align with each token’s semantic during the denoising time steps [15]. Based on this, we propose and compute the Relative Evolution Rate (RER) of the cross-attention maps for the <<EOS>> token relative to other tokens. The RER at each timestep is visualized in the heatmaps of Fig. 1, where red indicates a positive RER and blue indicates a negative RER. We observe that backdoor samples exhibit distinct feature evolution patterns at the <<EOS>> token compared to benign samples. To systematically quantify the dynamic anomalies at the <<EOS>> token, we propose two effective approaches.
In the first approach, we treat the attention maps corresponding to each token are spatially independent and introduce DAA-I. The dynamic feature of attention maps is measured using the Frobenius norm [39]. Furthermore, to better capture the interactions between attention maps, we propose a dynamical system-based approach, named DAA-S. Inspired by the complex network theory [40, 41, 42, 43, 44], we treat each attention map as a node and construct a graph that evolves over time. The model formulates the dynamic correlations among attention maps using a graph-based state equation and theoretically analyze the global asymptotic stability of this model. Since the state equation is an ordinary differential equation (ODE), we can numerically solve it to optimize the feature of each attention maps and calculate the corresponding dynamic feature. Both methods are independent of prompt length, achieving an effectiveness in a real-world application. Extensive experiments across six representative backdoor attack scenarios show the effectiveness of DAA. Our method achieves an average F1 score of 79.27% and an AUC of 86.27%, significantly outperforming the previous state-of-the-art results of 37.77% and 77.02% in terms of F1 Score and AUC, respectively.
@@
  • Experimental results show that our approach outperforms existing methods under six representative backdoor attack scenarios, achieving the average score of 79.27% F1 score and 86.27% AUC.
@@ II-B Backdoor Attack on T2I Diffusion Models
Inspired by backdoor attacks in image classification, recent works explore the vulnerabilities of text-to-image diffusion models to such attacks [23, 24, 25, 26, 27]. Struppek et al. introduce Rickrolling the Artist [23], a backdoor attack that embeds visually similar characters (e.g., Cyrillic “o”) into the text encoder. This method aligns the feature representations of backdoor and target prompts within the text embedding space while preserving the original feature embedding of benign samples. Chou et al. [25] propose Villan Diffusion, a framework that unifies backdoor attacks across both unconditional and conditional diffusion models. Their method incorporates trigger implantation into LoRA [31] and demonstrates its effectiveness across various training-free samplers. Zhai et al. introduce BadT2I [32], aiming to inject the backdoor through multimodal data poisoning. By leveraging a regularization loss, T2I diffusion models can be efficiently backdoored with only a few fine-tuning steps. To enable the precise backdoor customization, Wu et al. [27] leverage personalization techniques [24] to implant specific phrases as triggers. Specifically, they inject backdoors into textual inversion embeddings. Similarly, inspired by DreamBooth [16], Huang et al. [24] employ nouveau-tokens as triggers and align the predicted noise of backdoor samples with that of target samples. In contrast to prior approaches that require poisoned training data, Wang et al. [28] propose EvilEdit, a training-free and data-free backdoor attack. Inspired by model editing techniques [56, 57, 58], EvilEdit designs a closed-form way to modify projection matrices within cross-attention layers to implant backdoors. More recently, Zhang et al. introduce IBA [35], an approach designed to enhance the stealthiness of backdoor attacks. They analyze the anomalous traces left by existing backdoor samples, proposing syntactic triggers and a regularization loss term based on Kernel Mean Matching Distance (KMMD) [59] to bypass current detection.
Figure 2: The average relative evolution trajectories of the <<EOS>> token in benign samples (the orange line) and backdoor samples (the blue line) across six backdoor scenarios [23, 25, 28, 32, 24]. Each node represents the feature at time step $t$, where $t\in[0,50]$. The horizontal and vertical axes correspond to the first two principal components extracted using Principal Component Analysis (PCA) [60].
@@ II-C Backdoor Defense on T2I Diffusion Models
In response to the increasing security threat posed by backdoor attacks, several defense methods have been proposed. Wang et al. [33] propose T2IShield to detect, localize, and mitigate backdoor attacks. They find the “Assimilation Phenomenon” in backdoor samples, i.e., a consistent structural response in attention maps. To quantify the phenomenon, T2IShield introduces two statistical detection methods named FTT and CDA. Specifically, FTT utilizes the Frobenius norm [39] to quantify structural consistency, while CDA employs the covariance matrix to assess structural variations on the Riemannian manifold. Chew et al. [61] propose a text perturbation-based defense mechanism that applies character-level and word-level perturbations to backdoor samples. Although the method shows strong effect on preventing the backdoor from being triggered, they fail to effectively detect backdoor samples. Furthermore, Guan et al. [34] introduce UFID, a novel method that detects backdoor samples based on output diversity. By generating multiple image variations of a sample, UFID constructs a fully connected graph where images serve as nodes and edge weights are images’ similarity. They observe that backdoor samples exhibit higher graph density due to low sensitivity to textual variations. Beyond these text-conditioned defenses, several studies have also explored defenses in unconditional diffusion models [1, 62], aiming to detect backdoor samples [63], reverse backdoor triggers [64], or eliminate backdoor effects [65]. Although these works do not address textual triggers, they offer complementary perspectives for understanding and mitigating vulnerabilities in T2I diffusion models.
@@ III-A Preliminaries
where $T$ is the hyper-parameter for denoising time steps. For simplicity, we refer to “attention maps” as cross-attention maps generated by UNet [30] in the rest of the paper.
@@ III-B Anomalous Cue of Backdoor Samples
The tokens of a tokenized prompt $P$ are typically divided into three types: <<BOS>> token, prompt token, <<EOS>> token [29]. For example, as shown in Fig. 1, the input prompt is tokenized by “<<BOS>> a pen in the field <<EOS>>”, where <<BOS>> and <<EOS>> token serve as the beginning and ending signal to the prompt. Previous works have shown that <<EOS>> contains richer information than other tokens and plays a decisive role in the image generation [67, 68, 69]. In this work, we further identify the anomalous cue of backdoor samples is their distinct evolution pattern of the <<EOS>> token relative to other tokens. Formally, given a set of attention maps at time step $t$, i.e., $M^{t}=\{M^{t}_{1},M^{t}_{2},\dots,M^{t}_{L}\}$, we calculate Evolve Rate (ER) of each map by:
@@
where $t\in[0,T-1]$. Then, the Relative Evolve Rate (RER) of <<EOS>> token’s map is represented by:
@@
To better illustrate the distinct dynamic feature of $RER_{L}^{t}$ between benign samples and backdoor samples, we apply Principal Component Analysis (PCA) [60] to $RER_{L}^{t}$ and select first two dimension to visualization. Specifically, we compute the average RER of six backdoor attack scenarios mentioned in Table I, involving 6,700 backdoor samples and 6,700 benign samples to visualize the trajectories. As shown in Fig. 2, the two trajectories exhibit significant inconsistency. In particular, the trajectory points of the backdoor samples (i.e., the blue trajectory) are more densely clustered with smaller step sizes, indicating a slower state update. On the other hand, the benign samples (i.e., the yellow trajectory) have more sparsely distributed trajectory points with larger step sizes, suggesting a faster state update.
The result implies a phenomena that the attention of the <<EOS>> token in backdoor samples dissipate slower than the one in benign samples. Recall that the <<EOS>> plays a decisive role in generation process [67, 68, 69]. Intuitively, for a benign sample, <<EOS>> token dominates the early-stage diffusion process. Its attention gradually shifts from focusing on the <<EOS>> token to important tokens during the denoising process. This results in the <<EOS>> token having a higher rate of attention evolve compared to the average of other tokens. In contrast, for a backdoor sample, the model is forced to learn a specific pattern. The trigger interferes with this process by taking control of the global shape, thereby creating a competition between <<EOS>> token and the trigger token. As a result, the <<EOS>> token exhibits a lower rate of attention evolve compared to the average of other tokens.
(a) DAA-I.
(b) DAA-S.
Figure 4: The feature probability density visualization for 6,700 benign samples and 6,700 backdoor samples. (a) Feature probability density computed by DAA-I. (b) Feature probability density computed by DAA-S. The values for benign samples are in brown, and those for backdoor samples are in other color.
Based on the above observation, We introduce a novel backdoor detection framework named Dynamic Attention Analysis (DAA) to detect the anomalous dynamic attention feature of the <<EOS>> token on backdoor samples. The overview of our DAA is shown in Fig. 3. Specifically, given the tokenized prompt $P$, the model generates a set of cross-attention maps $\{M^{0},M^{1},\dots,M^{T-1},M^{T}\}$ during the denoising time steps. We propose two effective methods to quantify the dynamic features, i.e., DAA-I and DAA-S. DAA-I treats the tokens’ attention maps as spatially independent. DAA-S regards the maps as a graph and captures the dynamic features based on dynamical system.
@@ III-C Spatially Independent Approach
To model the anomalous dynamic pattern for the attention maps, we first assume that each map are spatially independent. Formally, given a set of attention maps at time step $t$, i.e., $M^{t}=\{M^{t}_{1},M^{t}_{2},\dots,M^{t}_{L}\}$, we calculate Evolve Rate (ER) in Eq. (4) by the Frobenius norm [39]:
[…]
@@
$\text{DAA-I}(P)=\mathbb{I}[\text{DAA-I}^{t}_{s}\leqslant\alpha_{I}].$ (8)
We visualize the distribution of $\text{DAA-I}^{3}_{2}$ for six types of backdoor attack scenarios in Fig. 4a, which involves a number of 6,700 backdoor samples and 6,700 benign samples. We observe a distribution shift between backdoor and benign samples. Besides, the DAA-I values for backdoor samples are mostly less than 0, whereas those for benign samples are greater than 0. This further confirms that the evolution rate of the <<EOS>> token in backdoor samples is lower than that of non-<<EOS>> tokens.